OpenStudy (anonymous):

4 x 5/7 x -3/-10 in simplest form

OpenStudy (anonymous):

Is this multiplying 3 fractions? Like:\[4 \times \frac{ 5 }{ 7 } \times \frac{ -3 }{ -10 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, start by multiplying all the numerators: 4 x 5 x -3 = and separately, the denominators: 7 x -10 =

OpenStudy (anonymous):

\[\frac{ 4 }{ 1 } \times \frac{ 5 }{ 7 } \times \frac{ -3 }{ -10 }= \frac{ 4*5*(-3) }{ 1*7*(-10) } = \frac{ -60 }{ -70 }\]